# OpenMMReasoner: Pushing the Frontiers for Multimodal Reasoning with an Open and General Recipe

### v1

@@ 4.4 Training Recipe
To identify the most suitable RL algorithm for multimodal reasoning, we compare GSPO [zheng2025groupsequencepolicyoptimization], DAPO [yu2025dapoopensourcellmreinforcement], and GRPO [shao2024deepseekmathpushinglimitsmathematical] under a unified setting. We evaluate their training dynamics in terms of stability, exploration, and efficiency, as illustrated in Fig.˜4. GSPO demonstrates faster convergence, higher rewards, and more stable behavior than DAPO and GRPO, achieving an effective balance between exploration and stability. While DAPO applies online filtering to remove zero-variance samples, it exhibits early entropy collapse and slower progress due to larger rollout requirements. GRPO shows moderate stability but converges more slowly. Based on these observations and the ablation results summarized in Sec.˜4.1, we adopt GSPO as the RL algorithm in our final training setup.
@@
To balance task accuracy and output formatting, we adopt a composite reward function similar to those in [dsr1_cite, openr1], which combines both aspects in a weighted manner. Specifically, the final reward $R$ for each sample is defined as:
@@
Sampling Strategies Although RL algorithms naturally produce stronger learning signals on more challenging problems, their training efficiency remains limited. Previous work [kimiteam2025kimik15scalingreinforcement] employs a difficulty-based sampler, starting with easy tasks and gradually shifting to harder ones. We first use Qwen3-VL-8B [qwen3vl2025] to track the pass rate of each problem as a difficulty metric and implement the following strategy: RL training begins with simpler tasks and progressively transitions to more complex ones. Ablation results in Sec.˜4.1 show that curriculum learning does not outperform the original mixed sampling approach; therefore, we adopt the no-sampling strategy in our final configuration.
@@
Our empirical findings show that while OpenVisionReasoner [wei2025openvisionreasonertransferring] achieves competitive performance, its response length becomes excessively long, raising concerns about reasoning efficiency. This observation highlights an important question: how can a model maintain strong reasoning performance while remaining computationally efficient and adaptable across different problem types. In our reinforcement learning implementation, we adopt a similar overlength penalty strategy as proposed in DAPO [yu2025dapoopensourcellmreinforcement] to mitigate the overthinking behavior and achieve a balanced trade-off between reasoning depth and efficiency. As shown in Fig.˜6, we evaluate models on two benchmarks, MMMU [yue2024mmmumassivemultidisciplinemultimodal] and We-Math [qiao2024wemathdoeslargemultimodal]. Our model demonstrates higher reasoning efficiency compared to OVR [wei2025openvisionreasonertransferring]. As shown in Fig.˜6, OVR requires excessively long reasoning trajectories to reach correct answers, whereas our model achieves a favorable balance between accuracy and efficiency, maintaining a reasonable reasoning budget while delivering superior overall performance.
@@ 4.5 Analysis and Insights
We present the evaluation results of our final models for both the SFT and RL stages in Tab.˜6. Building on the strong reasoning foundation established during the SFT stage, the RL phase further sharpens and stabilizes these capabilities, leading to improved and more consistent performance. After RL, our model achieves state-of-the-art results on benchmarks such as WeMath [qiao2024wemathdoeslargemultimodal], MathVerse [zhang2024mathversedoesmultimodalllm], and MathVista [lu2024mathvistaevaluatingmathematicalreasoning] and demonstrating consistent improvement compare to SFT.
@@
As the model’s overall reasoning ability improves through RL training, we observe the gradual emergence of textual reasoning behaviors, suggesting a transfer of reasoning competence from multimodal to purely linguistic domains. As shown in Fig.˜5, validation performance on AIME24, AIME25, and AMC23 steadily increases throughout training, reflecting continuous and measurable gains in text-based reasoning capabilities. For AIME24 and AIME25, the reported accuracy represents the average over eight rollout runs. Prior work [huan2025doesmathreasoningimprove] has demonstrated that enhancements in mathematical reasoning can positively transfer to other reasoning domains. Evaluation results in Tab.˜8 compare baseline and RL-trained models, showing that textual reasoning improves and strengthens in tandem with multimodal reasoning across all training stages. These findings extend previous observations, indicating that cross-domain reasoning skills acquired via multimodal RL can effectively transfer to purely textual tasks, further highlighting the shared cognitive foundations and underlying mechanisms across different modalities.
@@ 1.2 Evaluation Details
We now describe our evaluation setup for multimodal reasoning benchmarks. The SFT and RL models share the same evaluation configuration. We use the system prompt shown in  Tab.˜4 to ensure the model outputs both the reasoning trace and the final answer in an extractable format. The extracted answer is then validated using a two-stage process: a rule-based validator followed by an LLM-as-judge validator. We first apply the rule-based validator to minimize evaluation cost; if the answer cannot be verified, we fall back to the LLM-as-judge, using the prompt provided in  Tab.˜5. For all evaluations, we use a temperature of 0.0 for reproducibility and set the maximum generation length to 49K tokens, except for AIME, where we use a temperature of 1.0. We employ vLLM [kwon2023efficient] as the serving engine to accelerate inference.

### v2

@@ 4.4 Training Recipe
To identify the most suitable RL algorithm for multimodal reasoning, we compare GSPO [50], DAPO [44], and GRPO [30] under a unified setting. We evaluate their training dynamics in terms of stability, exploration, and efficiency, as illustrated in Fig.˜4. GSPO demonstrates faster convergence, higher rewards, and more stable behavior than DAPO and GRPO, achieving an effective balance between exploration and stability. While DAPO applies online filtering to remove zero-variance samples, it exhibits early entropy collapse and slower progress due to larger rollout requirements. GRPO shows moderate stability but converges more slowly. Based on these observations and the ablation results summarized in Sec.˜4.1, we adopt GSPO as the RL algorithm in our final training setup.
@@
To balance task accuracy and output formatting, we adopt a composite reward function similar to those in [10, 13], which combines both aspects in a weighted manner. Specifically, the final reward $R$ for each sample is defined as:
@@
Sampling Strategies Although RL algorithms naturally produce stronger learning signals on more challenging problems, their training efficiency remains limited. Previous work [33] employs a difficulty-based sampler, starting with easy tasks and gradually shifting to harder ones. We first use Qwen3-VL-8B [29] to track the pass rate of each problem as a difficulty metric and implement the following strategy: RL training begins with simpler tasks and progressively transitions to more complex ones. Ablation results in Sec.˜4.1 show that curriculum learning does not outperform the original mixed sampling approach; therefore, we adopt the no-sampling strategy in our final configuration.
@@
Our empirical findings show that while OpenVisionReasoner [37] achieves competitive performance, its response length becomes excessively long, raising concerns about reasoning efficiency. This observation highlights an important question: how can a model maintain strong reasoning performance while remaining computationally efficient and adaptable across different problem types. In our reinforcement learning implementation, we adopt a similar overlength penalty strategy as proposed in DAPO [44] to mitigate the overthinking behavior and achieve a balanced trade-off between reasoning depth and efficiency. As shown in Fig.˜6, we evaluate models on two benchmarks, MMMU [46] and We-Math [27]. Our model demonstrates higher reasoning efficiency compared to OVR [37]. As shown in Fig.˜6, OVR requires excessively long reasoning trajectories to reach correct answers, whereas our model achieves a favorable balance between accuracy and efficiency, maintaining a reasonable reasoning budget while delivering superior overall performance.
@@ 4.5 Analysis and Insights
We present the evaluation results of our final models for both the SFT and RL stages in Tab.˜6. Building on the strong reasoning foundation established during the SFT stage, the RL phase further sharpens and stabilizes these capabilities, leading to improved and more consistent performance. After RL, our model achieves state-of-the-art results on benchmarks such as WeMath [27], MathVerse [49], and MathVista [23] and demonstrating consistent improvement compare to SFT.
@@
As the model’s overall reasoning ability improves through RL training, we observe the gradual emergence of textual reasoning behaviors, suggesting a transfer of reasoning competence from multimodal to purely linguistic domains. As shown in Fig.˜5, validation performance on AIME24, AIME25, and AMC23 steadily increases throughout training, reflecting continuous and measurable gains in text-based reasoning capabilities. For AIME24 and AIME25, the reported accuracy represents the average over eight rollout runs. Prior work [12] has demonstrated that enhancements in mathematical reasoning can positively transfer to other reasoning domains. Evaluation results in Tab.˜8 compare baseline and RL-trained models, showing that textual reasoning improves and strengthens in tandem with multimodal reasoning across all training stages. These findings extend previous observations, indicating that cross-domain reasoning skills acquired via multimodal RL can effectively transfer to purely textual tasks, further highlighting the shared cognitive foundations and underlying mechanisms across different modalities.
@@ 1.2 Evaluation Details
We now describe our evaluation setup for multimodal reasoning benchmarks. The SFT and RL models share the same evaluation configuration. We use the system prompt shown in  Tab.˜4 to ensure the model outputs both the reasoning trace and the final answer in an extractable format. The extracted answer is then validated using a two-stage process: a rule-based validator followed by an LLM-as-judge validator. We first apply the rule-based validator to minimize evaluation cost; if the answer cannot be verified, we fall back to the LLM-as-judge, using the prompt provided in  Tab.˜5. For all evaluations, we use a temperature of 0.0 for reproducibility and set the maximum generation length to 49K tokens, except for AIME, where we use a temperature of 1.0. We employ vLLM [16] as the serving engine to accelerate inference.